OpenStudy (anonymous):

What is the solution set for |2x – 6| – 7 = 7?

OpenStudy (anonymous):

|2x - 6| = 14 that gives you 2 equations 2x-6 = 14 and 2x-6 = -14 solve for x for both of them 2x-6 = 14 ---> 2x = 20 --->x=10 2x-6 = -14 ---> 2x = -8 ---> x=-4 so your solution set is x = {-4,10}

OpenStudy (anonymous):

ok thank you soo much

OpenStudy (anonymous):

for positive charge we can write 2x-6-7=7 =>2x-13=7 =>2x=7+13 =>2x=20 =>x=10 for negative charge we can write -(2x-6)-7=7 =>-2x+6-7=7 =>-2x-1=7 =>-2x=7+1 =>-2x=8 =>x=-4